OpenStudy (anonymous):

Need help with one question please :) will give medal and fan. (**i tried graphing it but i think i did something wrong either way i think the answer is -1.5 or -2.5 but i am not sure**)

OpenStudy (anonymous):

@Alimara can you please help again?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

yes hi hello i can try to help

OpenStudy (anonymous):

Can I help

OpenStudy (anonymous):

yeah if you think you can help that'd be cool :)

OpenStudy (anonymous):

Yes, of course

OpenStudy (anonymous):

middle point of 2 points x1(−4,y1−2)x2(1,y23)(x2+x12,y2+y12)x2+x12⇐x-coordinate

OpenStudy (anonymous):

Take the average of the x-coordinates: (-4 + 1)/2 = -3/2 Take the average of the y-coordinates: (-2 + 3)/2 = 1/2 The midpoint is (-3/2, 1/2)

OpenStudy (anonymous):

Um I'm sorry I don't understand :/ >.<

OpenStudy (anonymous):

I am saying that the midpoint is (-3/2, 1/2)

OpenStudy (anonymous):

let me have a minute im going to try to make another graph

OpenStudy (anonymous):

1.5 most likely from what i can see

OpenStudy (anonymous):

that was my first choice yeah, giving you a medal but since i already fanned you i can't do that again lol